OpenStudy (anonymous):

40. If the expression x3 + 2hx - 2 is equal to 6 when x = -2, what is the value of h? A) 0 B) -2 C) -4 D) 4 E) 6

OpenStudy (anonymous):

I think its -4 I had a similar question

OpenStudy (anonymous):

idk if im right

OpenStudy (juanpablojr):

and you may be right lol

OpenStudy (here_to_help15):

c -4

OpenStudy (juanpablojr):

i knew it >_>

OpenStudy (owlcoffee):

if we have: \[x ^{3}+(2h)x-2=6\] And we are gicen the value of the variable, then all we have to do is replace the x there: \[(-2)^3 +(2h)(-2)-2=6\] ordering them up: \[-8-4h-2=6\] further: \[-4h-16=0\] And solving for "h": \[h=\frac{ 16 }{ (-4) }\] \[h=-4\]

OpenStudy (anonymous):

thanks for the explain!!! :)

OpenStudy (here_to_help15):

@PANTIEPOWER is this the next question? If -3/(a - 3) = 3/(a + 2), then a = ? A) -3 B) -2 C) 1/2 D) 2 E) 3